# Test-Time Alignment of LLMs via Sampling-Based Optimal Control in pre-logit space

### v1

@@ 5.2 Results
Table 1 lists average rewards, diversity score, and coherence score of each method. In terms of average reward, AISP achieves the highest among methods. AISP achieved up to about 40% improvement over BoN (top-$p$). AISP also outperforms RE-Control even though it does not require building training datasets. This result indicates that AISP is superior to baselines as a sampling-based reward optimization. ARGS does not work very well in our experiment. This is because ARGS needs to evaluate next token generation by a reward model. For this purpose, a token-level reward model $r(y_{t},\bm{y}_{<t})$ is more suitable than a trajectory-level reward model $r(\mathbf{x},\mathbf{y})$ used in our experiment. However, token-level reward models generally require additional training or specialized techniques (Yoon et al., 2024; Chakraborty et al., 2024).
@@
Table 2 lists the win rate for AISP vs BoN. To compute win rate, we sampled 100 pairs of prompts and responses at random, and GPT-4 judges whether the response from AISP or BoN is better. Since the values are averaged over three trials, they do not always sum to 100 %. This table shows that AISP has higher win rates than those of BoN (top-$p$) under almost all of conditions. The results of average rewards and win rates show that AISP aligns LLMs better than BoN: i.e., AISP can generate more helpful and harmless responses through maximization of rewards than BoN.
Table 2: Win rate for AISP vs BoN (top-$p$). Top: SHP and Bottom: HHRLHF.
@@
To compare sample efficiencies of AISP and BoN, Fig. 4 plots curves of reward values during iterations on SHP. In this figure, AISP (Mean at $k$) is $1/n\sum_{i}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$. AISP (Best at $k$) is $\max_{i}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$, and AISP (Best so far) is $\max_{i,1\leq j\leq k}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,j}))$, which is $r_{\mathrm{best}}$ in Algorithm 1 at $k$. BoN corresponds to $\max_{\bm{\mathrm{y}}\in\mathcal{Y}_{N}}r(\bm{\mathrm{x}},\bm{\mathrm{y}})$ using N=nkN\!=\!nk samples. These curves are also evaluated on randomly selected 100 pairs, and are averaged over data samples and three trials. This figure shows that though AISP underperforms BoN in the early iterations, it improves more rapidly and eventually surpasses BoN as the number of iterations increases. Additionally, while the maximum number of iterations $k$ was set to 32 in this experiment, it is likely that the performance gap would become more pronounced with a larger number of iterations. Reward values of AISP (Mean at $k$) and AISP (Best at $k$) also increase according to $k$. This indicates that AISP not only optimizes the resulting response but also optimizes the distribution of responses. Therefore, AISP obtains aligned distributions of response without any additional techniques. Note that additional experiments conducted on HHRLHF are presented in Appendix D.2, which show consistent trends with the main results.
@@
Figure 3: Sample efficiency to improve rewards: reward curve against $k$ iterations. For each iteration, both methods generate 32 samples.
@@
Table 3: KL divergence from the base LLM of AISP ($\lambda$, $\alpha$), ARGS, and RE-Control.
@@
5.3 Batched AISP
As above, AISP achieves higher rewards than BoN with fewer samples. However, it requires sequential process: $\kappa$ iterations. When we need to reduce time complexity, AISP can be accelerated by processing $b$ prompts with small $n$ in batches as discussed in Section 3.6. We compare Batched AISP with BoN of $N=128$ under the same iterations for processing multiple prompts; i.e., $\kappa=b$ and $N=nb$. In this experiment, we use Llama3_8B with UltraRM on 100 prompts in SHP and evaluate Batched AISP under multiple settings of $(b,n)$. In Fig. 4, $(b,n)$ of AISP1, AISP2, AISP3, and AISP4 correspond to (8,16), (16,8), (32,4), and (64,2), respectively. Fig. 4 demonstrates that AISP can outperform BoN even under the same iterations for $D$ prompts. In addition, this figure shows that AISP can exceed BoN if it has at least four samples per iteration.
5.4 KL divergence
@@ 6 Conclusion
In this paper, we propose adaptive importance sampling on a pre-logit distribution for alignment of LLMs. Our method assumes that pre-logit distributions are composed of a Gaussian perturbation and the pre-logit of the base LLM, and optimizes the Gaussian perturbation through importance sampling. Since our method is simple, future work could include combinations of AISP and fine-tuning, and different importance sampling techniques.
@@ Appendix A Proofs
$F(r,p,\bm{\mathrm{x}},\lambda)$ $=\mathrm{log}\left(\int\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right)\frac{q(V)}{q(V)}p(V)dV\right)$ (21)
@@
$F(r,p,\bm{\mathrm{x}},\lambda)$ $=\mathrm{log}\left(\mathbb{E}_{V\sim\mathbb{Q}}\left[\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right)\frac{p(V)}{q(V)}\right]\right)$ (24)
@@
$D_{\mathrm{KL}}(\mathbb{Q}|\mathbb{P})$ $=\int\log\left(\frac{q(V)}{p(V)}\right)q^{*}(V)dV$ (30)
@@
$w(V)$ $=\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))-\frac{1}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}_{t}-\frac{1}{2}\hat{\bm{u}}_{t}^{\top}\hat{\bm{u}}_{t}\right).$ (40)
@@
$=\operatorname*{arg\,min}_{U}\int q^{*}(V)\left(\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}(\bm{v}_{t}-\bm{u}_{t})\right)dV+\bm{u}_{t}^{\top}\bm{u}_{t}.$ (45)
@@
$\frac{1}{\partial\bm{u}_{t}}\left(\int q^{*}(V)\left(\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}(\bm{v}_{t}-\bm{u}_{t})\right)dV+\bm{u}_{t}^{\top}\bm{u}_{t}\right)=\int q^{*}(V)\bm{v}_{t}dV-\bm{u}^{\top},$ (46)
and thus, the optimal mean $U^{*}\!=\![\bm{u}^{*}_{1},\dots,\bm{u}^{*}_{\tau}]$ is obtained by
@@
$\textstyle\mathbb{E}_{V\sim\mathbb{Q}^{*}}[\bm{v}_{t}]=\int\bm{v}_{t}q^{*}(V)dV=\int\bm{v}_{t}\frac{q^{*}(V)}{q(V|\hat{U},\sigma^{2})}q(V|\hat{U},\sigma^{2})dV=\mathbb{E}_{V\sim\mathbb{Q}_{\hat{U},\sigma^{2}}}[w(V)\bm{v}_{t}],\$(48)
@@
$w(V)$ $=\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right)\frac{\frac{1}{(2\pi\sigma^{2})^{\frac{d\tau}{2}}}\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}\bm{v}_{t}\right)}{\frac{1}{(2\pi\sigma^{2})^{\frac{d\tau}{2}}}\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}(\bm{v}_{t}-\hat{\bm{u}}_{t})^{\top}(\bm{v}_{t}-\hat{\bm{u}}_{t})\right)}$ (49)
$=\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))-\frac{1}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}_{t}-\frac{1}{2}\hat{\bm{u}}_{t}^{\top}\hat{\bm{u}}_{t}\right),$ (50)
@@ Appendix B Algorithm
1:Hyper-parameters $\lambda$, $\alpha$, $\sigma^{2}$, $n$, and $\kappa$. reward models $r(\bm{\mathrm{x}},\bm{\mathrm{y}})$, Input prompt $\bm{\mathrm{x}}$
2:Initialization: $\hat{U}^{1}=\bm{O}$, $r_{\mathrm{best}}=$-$\infty$
3:for $k=1,\dots,\kappa$ do
4:  for $i=1,\dots,n$ do
5:   $V^{i,k}\sim q(V|\hat{U}^{k},\sigma^{2})$
6:   $\bm{y}^{i}_{<1}=\bm{\mathrm{x}}$ for $i=1,\dots,n$
7:   for $t=1,\dots T$ do
8:     We get $\bm{z}^{i}_{t}=\phi_{\mathrm{LLM}}(\bm{y}^{i}_{<t})$ by adding $\bm{y}^{i}_{<t}$ to LLM
9:     if $t\leq\tau$ then
10:      $\bm{z}^{i}_{t}=\bm{z}^{i}_{t}+\bm{v}^{i}_{t}$
@@
15:   Get rewards $r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$ by adding $\bm{\mathrm{y}}(V^{i,k})$ to the reward model
16:   if $r_{\mathrm{best}}<r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$ then
17:     $\bm{\mathrm{y}}_{\mathrm{best}}=\bm{\mathrm{y}}(V^{i,k})$ and $r_{\mathrm{best}}=r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$
18:  Compute weights $\bar{w}^{i}$ by Eq. (19) for $i=1,\dots,n$
19:  $\hat{U}^{k+1}=[\hat{\bm{u}}^{k+1}_{1},\dots,\hat{\bm{u}}^{k+1}_{\tau}]$ by $\hat{\bm{u}}^{k+1}_{t}=\sum_{i}\bar{w}^{i}\bm{v}^{i,k}_{t}$
20:We generate $\bm{\mathrm{y}}(U^{*})\!=\!\mathrm{arg}\!\max_{V\in\mathcal{V}}\bm{\mathrm{y}}(V)$ where $\mathcal{V}=\{V^{i}|V^{i}\sim q(V|\hat{U}^{\kappa},\sigma^{2})\}$
21:if $r_{\mathrm{best}}<r(\bm{\mathrm{x}},\bm{\mathrm{y}}(U^{*}))$ then
22:  $\bm{\mathrm{y}}_{\mathrm{best}}=\bm{\mathrm{y}}(U^{*})$ and $r_{\mathrm{best}}=r(\bm{\mathrm{x}},\bm{\mathrm{y}}(U^{*}))$
23:Return $\bm{\mathrm{y}}_{\mathrm{best}}$ and $r_{\mathrm{best}}$
Algorithm 1 is the pseudo code of AISP. First, we generate $V^{i}$ from the prior distribution in Line 6 and generate responses $y(V^{i})$ in Lines 7-14. Line 11 decodes a token based on pre-logit. Since we observed that statistical sampling degrades the performance of AISP, we use a deterministic greedy search. Next, we evaluate reward values for each $\bm{\mathrm{y}}(V^{i})$ in Line 15. Line 16 stores the best response during AISP because we select the best $\bm{\mathrm{y}}$ among $n\kappa$ samples as the results like BoN. After reward evaluation, we update $\hat{U}$ in Lien 19. Finally, we generate $\bm{\mathrm{y}}(U^{*})$ as a last candidate of response and compare it with $\bm{\mathrm{y}}_{\mathrm{best}}$. Note that though adaptive importance sampling generally uses $n\kappa$ samples, i.e., all generating samples, at the last iteration, we only use the $n$ generated samples for each iteration due to computational cost. If we need multiple responses, AISP can be modified to output the set of $\bm{\mathrm{y}}$ by using top-$k$ in Lines 17 and 22.

### v2

@@ 5.2 Results
Table 1 lists average rewards, diversity score, and coherence score of each method. In terms of average reward, AISP achieves the highest among methods. AISP achieved up to about 40% improvement over BoN (top-$p$). AISP also outperforms RE-Control even though it does not require building training datasets. This result indicates that AISP is superior to baselines as a sampling-based reward optimization. ARGS does not work very well in our experiment. This is because ARGS needs to evaluate next token generation by a reward model. For this purpose, a token-level reward model $r(y_{t},\bm{y}_{<t})$ is more suitable than a trajectory-level reward model $r(\mathbf{x},\mathbf{y})$ used in our experiment. However, token-level reward models generally require additional training or specialized techniques (Yoon et al., 2024; Chakraborty et al., 2024; Xu et al., 2025). Additionally, BoN with gaussian pre-logit noise (BoN ($\mathcal{N}$)) achieves competitive performance with BoN (top-$p$). This implies that the Gaussian assumption for pre-logits is not very strong, and it generates good candidate responses even without importance sampling.
@@
Table 2 lists the win rate for AISP vs BoN (top-$p$). To compute win rate, we sampled 100 pairs of prompts and responses at random, and GPT-4 judges whether the response from AISP or BoN is better. Since the values are averaged over three trials, they do not always sum to 100 %. This table shows that AISP has higher win rates than those of BoN (top-$p$) under almost all of conditions. The results of average rewards and win rates show that AISP aligns LLMs better than BoN: i.e., AISP can generate more helpful and harmless responses through maximization of rewards than BoN.
Table 2: Win rate for AISP vs BoN (top-$p$) on SHP (top) and HHRLHF (bottom): (L: Llama3_8B, V: Vicuna_7B, G: Gemma3_4B, U: UltraRM, E: Eurus)
L & U L & E V & U V & E G & U G & E
@@
To compare sample efficiencies of AISP and BoN, Fig. 3 plots curves of reward values during iterations on SHP. In this figure, AISP (Mean at $k$) is $1/n\sum_{i}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$. AISP (Best at $k$) is $\max_{i}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$, and AISP (Best so far) is $\max_{i,1\leq j\leq k}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,j}))$, which is $r_{\mathrm{best}}$ in Algorithm 1 at $k$. BoN corresponds to $\max_{\bm{\mathrm{y}}\in\mathcal{Y}_{N}}r(\bm{\mathrm{x}},\bm{\mathrm{y}})$ using N=nkN\!=\!nk samples. These curves are also evaluated on randomly selected 100 pairs, and are averaged over data samples and three trials. This figure shows that though AISP underperforms BoN in the early iterations, it improves more rapidly and eventually surpasses BoN as the number of iterations increases. In addition, while the maximum number of iterations $k$ was set to 32 in this experiment, it is likely that the performance gap would become more pronounced with a larger number of iterations. Reward values of AISP (Mean at $k$) and AISP (Best at $k$) also increase according to $k$. This indicates that AISP not only optimizes the resulting response but also optimizes the distribution of responses. Thus, AISP obtains aligned distributions of response without any additional techniques. Note that additional experiments conducted on HHRLHF are presented in Appendix D.2, which show consistent trends with the main results.
@@
Figure 3: Sample efficiency comparison: Reward curves over iterations. Both methods generate 32 samples per iteration.
@@
5.3 Batched AISP
As above, AISP achieves higher rewards than BoN with fewer samples. However, it requires sequential process: $\kappa$ iterations. When we need to reduce time complexity, AISP can be accelerated by processing $b$ prompts with small $n$ in batches as discussed in Section 3.6. We compare Batched AISP with BoN of $N=128$ under the same iterations for processing multiple prompts; i.e., $\kappa\!=\!b$ and N=nbN\!=\!nb. In this experiment, we use Llama3_8B with UltraRM on 100 prompts in SHP and evaluate Batched AISP under multiple settings of $(b,n)$. In Fig. 4, $(b,n)$ of AISP1, AISP2, AISP3, and AISP4 correspond to (8,16), (16,8), (32,4), and (64,2), respectively. Fig. 4 demonstrates that AISP can outperform BoN even under the same iterations for $D$ prompts. In addition, this figure shows that AISP can exceed BoN if it has at least four samples per iteration.
Runtime evaluation
To investigate the computation cost of AISP, we evaluate the runtime of AISP on a stand-alone server (A100 VRAM 40GB). For fair comparison, we set the generated token length to fixed 128. First, to confirm that the overhead for weight updating Eq. (14) is trivial as explained in Section 3.6, we evaluate the runtimes for one iteration of AISP (n=32n\!=\!32) and BoN (N=32N\!=\!32) averaged over 10 prompts in SHP with Llama3_8B and UltraRM. AISP and BoN take 7.75 s and 7.68 s, respectively. Thus, the overhead for weight updating is only 1%. Next, we evalauted the runtime for Batched AISP (b,n)=(4,8)(b,n)\!=\!(4,8) and BoN (N=32N\!=\!32) on 100 prompts in SHP with the same models. Batched AISP and BoN take 9.35 s and 6.83 s per prompt, respectively. The overhead reaches 36.8 % because the runtime of Batched AISP for one mini-batch is determined by the largest prompts in the mini-batch. After sorting the prompts in terms of the token length, the runtime of Batched AISP becomes 7.38 s/prompt, and thus, the overhead is just 8 %.
5.4 KL divergence
Table 3: KL divergence from the base LLM.
@@ 6 Conclusion
In this paper, we propose adaptive importance sampling on a pre-logit distribution for alignment of LLMs. Our method injects a Gaussian perturbation to the pre-logit of the base LLM, and optimizes the Gaussian perturbation through importance sampling. Since our method is simple, future work could include combinations of AISP and fine-tuning, and different importance sampling techniques.
Impact Statement
This paper presents work whose goal is to advance the field of LLM and its applications. Malicious users may be able to increase the aggressiveness of LLMs by using reward models for inappropriate purposes. However, we believe that the risk of such misuse is common across machine learning technologies in general, and one of our proposal is not particularly significant.
@@ Appendix A Proofs
$F(r,p,\bm{\mathrm{x}},\lambda)$
$=$ $\mathrm{log}\left(\int\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right)\frac{q(V)}{q(V)}p(V)dV\right)$ (20)
@@
$F(r,p,\bm{\mathrm{x}},\lambda)$
@@
$D_{\mathrm{KL}}(\mathbb{Q}|\mathbb{P})$
$=$ $\int\log\left(\frac{q(V)}{p(V)}\right)q^{*}(V)dV$ (28)
@@
$w(V)$ $=\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))-\frac{1}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}_{t}-\frac{1}{2\sigma^{2}}\hat{\bm{u}}_{t}^{\top}\hat{\bm{u}}_{t}\right).$ (38)
@@
$=$ $\operatorname*{arg\,min}_{U}\int q^{*}(V)\left(\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}(\bm{v}_{t}-2\bm{u}_{t})\right)dV+\sum_{t=1}^{\tau}\bm{u}_{t}^{\top}\bm{u}_{t}.$ (43)
@@
$\frac{1}{\partial\bm{u}_{t}}\left(\int q^{*}(V)\left(\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}(\bm{v}_{t}-2\bm{u}_{t})\right)dV+\sum_{t=1}^{\tau}\bm{u}_{t}^{\top}\bm{u}_{t}\right)$
$=-2\int q^{*}(V)\bm{v}_{t}dV+2\bm{u}_{t}.$ (44)
Thus, the optimal mean $U^{*}\!=\![\bm{u}^{*}_{1},\dots,\bm{u}^{*}_{\tau}]$ is obtained by
@@
$\textstyle\mathbb{E}_{V\sim\mathbb{Q}^{*}}[\bm{v}_{t}]$ $=\int\bm{v}_{t}q^{*}(V)dV$ (46)
$=\int\bm{v}_{t}\frac{q^{*}(V)}{q(V|\hat{U},\sigma^{2})}q(V|\hat{U},\sigma^{2})dV$ (47)
$=\mathbb{E}_{V\sim\mathbb{Q}_{\hat{U},\sigma^{2}}}[w(V)\bm{v}_{t}],\$(48)
@@
$w(V)=\textstyle\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right)\frac{\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}\bm{v}_{t}\right)}{\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}(\bm{v}_{t}-\hat{\bm{u}}_{t})^{\top}(\bm{v}_{t}-\hat{\bm{u}}_{t})\right)}$
$=$ $\textstyle\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))-\frac{1}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}_{t}-\frac{1}{2\sigma^{2}}\hat{\bm{u}}_{t}^{\top}\hat{\bm{u}}_{t}\right),$ (49)
@@ Appendix B Algorithm
0: Hyper-parameters $\lambda$, $\alpha$, $\sigma^{2}$, $n$, and $\kappa$. reward models $r(\bm{\mathrm{x}},\bm{\mathrm{y}})$, Input prompt $\bm{\mathrm{x}}$
1: Initialization: $\hat{U}^{1}=\bm{O}$, $r_{\mathrm{best}}=$-$\infty$
2:for $k=1,\dots,\kappa$ do
3:  for $i=1,\dots,n$ do
4:   $V^{i,k}\sim q(V|\hat{U}^{k},\sigma^{2})$
5:   $\bm{y}^{i}_{<1}=\bm{\mathrm{x}}$ for $i=1,\dots,n$
6:   for $t=1,\dots T$ do
7:    We get $\bm{z}^{i}_{t}=\phi_{\mathrm{LLM}}(\bm{y}^{i}_{<t})$ by adding $\bm{y}^{i}_{<t}$ to LLM
8:    if $t\leq\tau$ then
9:     $\bm{z}^{i}_{t}=\bm{z}^{i}_{t}+\bm{v}^{i}_{t}$
10:    end if
@@
15:    end if
16:   end for
17:   Get rewards $r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$ by adding $\bm{\mathrm{y}}(V^{i,k})$ to the reward model
18:   if $r_{\mathrm{best}}<r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$ then
19:    $\bm{\mathrm{y}}_{\mathrm{best}}=\bm{\mathrm{y}}(V^{i,k})$ and $r_{\mathrm{best}}=r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i,k}))$
20:   end if
21:  end for
22:  Compute weights $\bar{w}^{i}$ by Eq. (18) for $i=1,\dots,n$
23:  $\hat{U}^{k+1}=[\hat{\bm{u}}^{k+1}_{1},\dots,\hat{\bm{u}}^{k+1}_{\tau}]$ by $\hat{\bm{u}}^{k+1}_{t}=\sum_{i}\bar{w}^{i}\bm{v}^{i,k}_{t}$
24:end for
25: We generate $\bm{\mathrm{y}}(U^{*})\!=\!\mathrm{arg}\!\max_{V\in\mathcal{V}}\bm{\mathrm{y}}(V)$ where $\mathcal{V}=\{V^{i}|V^{i}\sim q(V|\hat{U}^{\kappa},\sigma^{2})\}$
26:if $r_{\mathrm{best}}<r(\bm{\mathrm{x}},\bm{\mathrm{y}}(U^{*}))$ then
27:  $\bm{\mathrm{y}}_{\mathrm{best}}=\bm{\mathrm{y}}(U^{*})$ and $r_{\mathrm{best}}=r(\bm{\mathrm{x}},\bm{\mathrm{y}}(U^{*}))$
28:end if
29:Return $\bm{\mathrm{y}}_{\mathrm{best}}$ and $r_{\mathrm{best}}$
Algorithm 1 is the pseudo code of AISP. First, we generate $V^{i}$ from the prior distribution in Line 5 and generate responses $y(V^{i})$ in Lines 6-16. Line 11 decodes a token based on pre-logit. Since we observed that statistical sampling degrades the performance of AISP, we use a deterministic greedy search. Next, we evaluate reward values for each $\bm{\mathrm{y}}(V^{i})$ in Line 17. Line 18 stores the best response during AISP because we select the best $\bm{\mathrm{y}}$ among $n\kappa$ samples as the results like BoN. After reward evaluation, we update $\hat{U}$ in Line 23. Finally, we generate $\bm{\mathrm{y}}(U^{*})$ as a last candidate of response and compare it with $\bm{\mathrm{y}}_{\mathrm{best}}$. Note that though adaptive importance sampling generally uses $n\kappa$ samples, i.e., all generating samples, at the last iteration, we only use the $n$ generated samples for each iteration due to computational cost. If we need multiple responses, AISP can be modified to output the set of $\bm{\mathrm{y}}$ by using top-$k$ in Lines 19 and 27.